Jerielle Pelley is the front office manager at the 125-room limited-service Best Stay Inn. She also...

  1. Jerielle Pelley is the front office manager at the 125-room limited-service Best Stay Inn. She also serves as her property’s RM. Jerielle has just taken a call from Lawrence, a friend and the RM at a hotel within her comp set. Because of an internal oversight, Lawrence’s hotel is overbooked by 70 group rooms next Saturday. Lawrence would like to purchase that number of rooms from Jerielle at their previously agreed upon walk rate of $75.00. Jerielle’s normal rack rate is $129.00. Currently, she has 55 occupied rooms (arrivals and stayovers) on the books for that day. She estimates that she could sell, at her normal ADR, another 30 rooms by Saturday. Complete the source of business worksheet and then answer the questions that follow.

Group

With Lawrence Walks

W/O Lawrence Walks

Sold

ADR

$129

Total rev estimate

Daily per room ancillary revenue

$8

$8

Total ancillary rev
Answer #1
Group With Lawrence Walks W/O Lawrence Walks
Sold 125 85
ADR $                              98.76 $                           129.00
Total rev estimate $                      12,345.00 $                     10,965.00
Daily per room ancillary revenue 8 8
Total ancillary rev $                        1,000.00 $                           680.00

Working:

Group With Lawrence Walks W/O Lawrence Walks
Sold =55+70 =55+30
ADR =12345/125 129
Total rev estimate =(55*129)+(70*75) =85*129
Daily per room ancillary revenue 8 8
Total ancillary rev =125*8 =85*8